My teacher gave me a problem to solve and we've been doing these kind off equations all week and I just don't get them....The growth rate of the world is 2.67% per year.If the present population of earth is 8 billion people, how long before the population doubles?
Answer by stanbon(75887)

The growth rate of the world is 2.67% per year.If the present population of earth is 8 billion people, how long before the population doubles?
-------------------
A(t) = ab^x
16 = 8(1+0.0267)^x
---
1.0267^x = 2
---
x(log(1.0267)) = log(2)
---
x = log(2)/log(1.0267)
---
x = 26.31 years
=================
Cheers,
Stan H.